Given that P=25-50Q
TC=100-15Q+60Q2
Where p=price,Q-quantity and TC-Total cost Get the level of output that minimises cosy
1
Expert's answer
2018-10-08T14:51:08-0400
P = 25 - 50Q
TC = 100 - 15Q + 60Q^2
The level of output that minimises cost is at TC' = MC = 0, so:
(100 - 15Q + 60Q^2)' = 0,
120Q - 15 = 0,
Q = 15/120 = 0.125.